# Lense-Thirring Acoustic Black Holes : Shadows and Light

### v1

@@ 6.2 Graphics and results
In this part of the paper, we use the obtained results and apply them to our case which correspond to the acoustic black hole in the slowly rotating regime. Indeed, with the use of equations (3.24), (6.11) and (6.12), we find
$\Omega_{\theta}$ $=\frac{aM\sin(\theta)\left(\xi(r-6M)(r-2M)+r^{2}\right)}{r(2M-r)\left(r^{2}-2M\xi(r-2M)\right)\sqrt{\Gamma}},$ (6.13)
$\Omega_{r}$ $=\frac{aM\sin(2\theta)\left(\xi(r-2M)^{2}+r^{2}\right)}{r^{3}\sin(\theta)\sqrt{\Gamma}},$ (6.14)
@@
$\Gamma=r^{4}-\frac{4a^{2}M^{2}\sin^{2}(\theta)\left(\xi(r-2M)^{2}+r^{2}\right)^{2}}{r(2M-r)\left(r^{2}-2M\xi(r-2M)\right)}.$ (6.15)
@@
$\left.\left.+2\xi\,r^{4}(2M-r)\left(24M^{2}-14Mr+3r^{2}\right)+r^{6}(8M-3r)\right)\right]^{\frac{1}{2}},$ (6.16)
@@
$\Omega=\sqrt{-\frac{a^{2}M^{2}(\cos(2\theta)(8M-3r)+8M-5r)}{2\left(r^{2}(2M-r)\left(2a^{2}M^{2}(\cos(2\theta)-1)+r^{3}(2M-r)\right)\right)}}.$ (6.17)
This result matches perfectly the Kerr magnitude after some trigonometric rearrangements [87]. To reveal the behavior of such a quantity, we illustrate in figure (6) its behavior as a function of $r$ for different values of the acoustic parameter $\xi$ and the rotation $a$.
@@
$r_{{\Omega=0}}^{-}$ $=\frac{2M\left(2\xi-\sqrt{(\xi-3)\xi}\right)}{\xi+1},$ (6.18)
$r_{{\Omega=0}}^{+}$ $=\frac{2M\left(2\xi+\sqrt{(\xi-3)\xi}\right)}{\xi+1}.$ (6.19)
From the expression of the $\Omega$ given in equation (6.16), it can be seen that the angular velocity is null if $a=0$. Such a case suggests that the Lense-Thirring effect is null for non rotating black hole. In our case, we have found that similar scenarios exist in the LTABH for values of the spin different than zero. Indeed, equations (6.18)-(6.19) show regions where the spacetime is not dragged by the black hole rotation parameter. This means that the particles orbits would not be subject of the precession occurring due to the frame dragging effect. Such a result is essentially due to the acoustic parameter that cancels the rotating effect.
@@ 7 Deflection angle
The metric (3.24) describes a slowly rotating acoustic black hole. We begin by rewriting such a metric in the equatorial plan in order to get the required the deflection angle of light. Indeed for $\theta=\frac{\pi}{2}$, we have
$ds^{2}=-\left[F(r)+2a\left(1-F(r)\right)\frac{d\phi}{dt}\right]dt^{2}+\frac{dr^{2}}{F(r)}+r^{2}d\phi^{2}.$ (7.1)
The equations given in (LABEL:dtdtau) and (5.7) can be used to compute the factor $\frac{d\phi}{dt}$ in this equation. The expression for such a quantity as a function of the impact parameter $b=\frac{E}{L}$, can be given by
$\frac{d\phi}{dt}=\frac{F(r)+ba\left(1-F(r)\right)}{br^{2}-a\left(1-F(r)\right)}.$ (7.2)
@@
$dr_{*}$ $=\frac{dr}{\sqrt{F(r)\left(F(r)+2a\left(1-F(r)\right)\frac{d\phi}{dt}\right)}},$ (7.3)
$f(r_{*})$ $=\frac{r}{\sqrt{F(r)+2a\left(1-F(r)\right)\frac{d\phi}{dt}}},$ (7.4)
@@
$\iint_{D_{R}}K_{\mathcal{G}}dS+\oint_{\partial D_{R}}kdt+\sum n_{i}=2\pi\chi\left(D_{R}\right).$ (7.6)
@@
$k\left(\gamma_{R}\right)=\big|\nabla_{\gamma_{R}}\dot{\gamma_{R}}\big|.$ (7.7)
Given that $\gamma_{R}=R=cte$ is verified by the geodesic $\gamma_{R}$ [91], where $R$ is the ”geometrical size” of the region $D_{R}$, the radial component of $k\left(\gamma_{R}\right)$ becomes
$\left(\nabla_{\gamma_{R}}\dot{\gamma_{R}}\right)^{r}=\dot{\gamma_{R}}^{\phi}+\partial_{\phi}\dot{\gamma_{R}}^{r}+\Gamma^{r}_{\phi\phi}\left(\dot{\gamma_{R}}^{\phi}\right)^{2}.$ (7.8)
Following [91], the second term provides
$\oint_{\partial D_{R}}kdt=\pi+\widehat{\alpha}$. (7.9)
@@
$\widehat{\alpha}=-\int_{0}^{\pi}\int_{\frac{b}{\sin\phi}}^{\infty}K_{\mathcal{G}}dS,$ (7.10)
@@
$K_{\mathcal{G}}=\frac{\mathbf{R}}{2},$ (7.11)
@@
$K_{\mathcal{G}}=\frac{18abM(\xi+1)}{r^{5}}-\frac{2M(\xi+1)}{r^{3}}.$ (7.12)
@@
$\widehat{\alpha}=\frac{4M(\xi+1)}{b}-\frac{4aM(\xi+1)}{b^{2}}+\mathcal{O}\left(M^{2},a^{2}\right),$ (7.13)
where $M$ higher orders are not included. At the limit $\xi\to 0$, the deflection angle (7.13) becomes
$\widehat{\alpha}\simeq\frac{4M}{b}\mp\frac{4aM}{b^{2}},$ (7.14)
where the $\mp$-branches would pertain to the pro-/retrograde motion. The obtained result matches precisely the LTBH black hole deflection of light [92]. To analyse the deflection of light by a slowly rotating acoustic black hole, we plot the associated behaviors as a function of the impact parameter $b$ in figure (7) for different values of the acoustic parameter $\xi$ and the spin $a$.
@@ Appendix A On the placement of critical radii
Further investigation could be carried regarding the LTABH who’s spacetime is described by the metric of equation (3.24). Indeed, we have seen that such a spacetime strongly depend on the value of the acoustic parameter $\xi$. While the horizon radius is always equal to $2M$, distinctions are also present. For convenience, we use the following notations
$r_{1}:=r_{ps},\,\,R_{s}\left(r_{1}\right)=R_{s}\left(r_{ps}\right):=R_{s},\quad r_{2}:=r_{as},\,\,R_{s}\left(r_{2}\right)=R_{s}\left(r_{as}\right):=R_{as}.$ (A.1)
@@ Appendix B The matrix $\mathcal{G}_{\mu\nu}$
$\mathcal{G}_{\mu\nu}=\begin{pmatrix}\mathcal{G}_{tt}&\mathcal{G}_{tr}&\mathcal{G}_{t\theta}&\mathcal{G}_{t\phi}\\ \mathcal{G}_{rt}&\mathcal{G}_{rr}&\mathcal{G}_{r\theta}&\mathcal{G}_{r\phi}\\ \mathcal{G}_{\theta t}&\mathcal{G}_{\theta r}&\mathcal{G}_{\theta\theta}&\mathcal{G}_{\theta\phi}\\ \mathcal{G}_{\phi t}&\mathcal{G}_{\phi r}&\mathcal{G}_{\phi\theta}&\mathcal{G}_{\phi\phi}\end{pmatrix}$, (B.1)
@@
$\mathcal{G}_{tt}$ $=\frac{c_{s}}{\sqrt{c_{s}^{2}-v_{\alpha}v^{\alpha}}}\,g_{tt}\left(c_{s}^{2}-v_{i}v^{i}\right),$ (B.2)
$\mathcal{G}_{tr}$ $=\frac{c_{s}}{\sqrt{c_{s}^{2}-v_{\alpha}v^{\alpha}}}\,\left(-v_{t}v_{r}+g_{tr}(c_{s}^{2}-v_{i}v^{i})\right),$ (B.3)
$\mathcal{G}_{t\theta}$ $=\frac{c_{s}}{\sqrt{c_{s}^{2}-v_{\alpha}v^{\alpha}}}\,\left(-v_{t}v_{\theta}+g_{t\theta}(c_{s}^{2}-v_{i}v^{i})\right),$ (B.4)
$\mathcal{G}_{t\phi}$ $=\frac{c_{s}}{\sqrt{c_{s}^{2}-v_{\alpha}v^{\alpha}}}\,\left(-v_{t}v_{\phi}+g_{t\phi}(c_{s}^{2}-v_{i}v^{i})\right),$ (B.5)
$\mathcal{G}_{rt}$ $=\frac{c_{s}}{\sqrt{c_{s}^{2}-v_{\alpha}v^{\alpha}}}\,\left(-v_{r}v_{t}+g_{rt}(c_{s}^{2}-v_{t}v^{t})\right),$ (B.6)
$\mathcal{G}_{rr}$ $=\frac{c_{s}}{\sqrt{c_{s}^{2}-v_{\alpha}v^{\alpha}}}\,\left(g_{rr}(c_{s}^{2}-v_{\alpha}v^{\alpha})+v_{r}v_{r}\right),$ (B.7)
$\mathcal{G}_{r\theta}$ $=\frac{c_{s}}{\sqrt{c_{s}^{2}-v_{\alpha}v^{\alpha}}}\,\left(g_{r\theta}(c_{s}^{2}-v_{\alpha}v^{\alpha})+v_{r}v_{\theta}\right),$ (B.8)
$\mathcal{G}_{r\phi}$ $=\frac{c_{s}}{\sqrt{c_{s}^{2}-v_{\alpha}v^{\alpha}}}\,\left(g_{r\phi}(c_{s}^{2}-v_{\alpha}v^{\alpha})+v_{r}v_{\phi}\right),$ (B.9)
$\mathcal{G}_{\theta t}$ $=\frac{c_{s}}{\sqrt{c_{s}^{2}-v_{\alpha}v^{\alpha}}}\,\left(-v_{\theta}v_{t}+g_{\theta t}(c_{s}^{2}-v_{t}v^{t})\right),$ (B.10)
$\mathcal{G}_{\theta r}$ $=\frac{c_{s}}{\sqrt{c_{s}^{2}-v_{\alpha}v^{\alpha}}}\,\left(g_{\theta r}(c_{s}^{2}-v_{\alpha}v^{\alpha})+v_{\theta}v_{r}\right),$ (B.11)
$\mathcal{G}_{\theta\theta}$ $=\frac{c_{s}}{\sqrt{c_{s}^{2}-v_{\alpha}v^{\alpha}}}\,\left(g_{\theta\theta}(c_{s}^{2}-v_{\alpha}v^{\alpha})+v_{\theta}v_{\theta}\right),$ (B.12)
$\mathcal{G}_{\theta\phi}$ $=\frac{c_{s}}{\sqrt{c_{s}^{2}-v_{\alpha}v^{\alpha}}}\,\left(g_{\theta\phi}(c_{s}^{2}-v_{\alpha}v^{\alpha})+v_{\theta}v_{\phi}\right),$ (B.13)
$\mathcal{G}_{\phi t}$ $=\frac{c_{s}}{\sqrt{c_{s}^{2}-v_{\alpha}v^{\alpha}}}\,\left(-v_{\phi}v_{t}+g_{\phi t}(c_{s}^{2}-v_{t}v^{t})\right),$ (B.14)
$\mathcal{G}_{\phi r}$ $=\frac{c_{s}}{\sqrt{c_{s}^{2}-v_{\alpha}v^{\alpha}}}\,\left(g_{\phi r}(c_{s}^{2}-v_{\alpha}v^{\alpha})+v_{\phi}v_{r}\right),$ (B.15)
$\mathcal{G}_{\phi\theta}$ $=\frac{c_{s}}{\sqrt{c_{s}^{2}-v_{\alpha}v^{\alpha}}}\,\left(g_{\phi\theta}(c_{s}^{2}-v_{\alpha}v^{\alpha})+v_{\phi}v_{\theta}\right),$ (B.16)
$\mathcal{G}_{\phi\phi}$ $=\frac{c_{s}}{\sqrt{c_{s}^{2}-v_{\alpha}v^{\alpha}}}\,\left(g_{\phi\phi}(c_{s}^{2}-v_{\alpha}v^{\alpha})+v_{\phi}v_{\phi}\right).$ (B.17)

### v2

@@ 6.2 Graphics and results
In this part of the paper, we use the obtained results and apply them to our case which correspond to the acoustic black hole in the slowly rotating regime. Indeed, with the use of equations (24), (70) and (71), we find
$\Omega_{\theta}$ $=\frac{aM\sin(\theta)\left(\xi(r-6M)(r-2M)+r^{2}\right)}{r(2M-r)\left(r^{2}-2M\xi(r-2M)\right)\sqrt{\Gamma}},$ (72)
$\Omega_{r}$ $=\frac{aM\sin(2\theta)\left(\xi(r-2M)^{2}+r^{2}\right)}{r^{3}\sin(\theta)\sqrt{\Gamma}},$ (73)
@@
$\Gamma=r^{4}-\frac{4a^{2}M^{2}\sin^{2}(\theta)\left(\xi(r-2M)^{2}+r^{2}\right)^{2}}{r(2M-r)\left(r^{2}-2M\xi(r-2M)\right)}.$ (74)
@@
$\left.\left.+2\xi\,r^{4}(2M-r)\left(24M^{2}-14Mr+3r^{2}\right)+r^{6}(8M-3r)\right)\right]^{\frac{1}{2}},$ (75)
@@
$\Omega=\sqrt{-\frac{a^{2}M^{2}(\cos(2\theta)(8M-3r)+8M-5r)}{2\left(r^{2}(2M-r)\left(2a^{2}M^{2}(\cos(2\theta)-1)+r^{3}(2M-r)\right)\right)}}.$ (76)
This result matches perfectly the Kerr magnitude after some trigonometric rearrangements bodyT . To reveal the behavior of such a quantity, we illustrate in figure (6) its behavior as a function of $r$ for different values of the acoustic parameter $\xi$ and the rotation $a$.
@@
$r_{{\Omega=0}}^{-}$ $=\frac{2M\left(2\xi-\sqrt{(\xi-3)\xi}\right)}{\xi+1},$ (77)
$r_{{\Omega=0}}^{+}$ $=\frac{2M\left(2\xi+\sqrt{(\xi-3)\xi}\right)}{\xi+1}.$ (78)
From the expression of the $\Omega$ given in equation (75), it can be seen that the angular velocity is null if $a=0$. Such a case suggests that the Lense-Thirring effect is null for non rotating black hole. In our case, we have found that similar scenarios exist in the LTABH for values of the spin different than zero. Indeed, equations (77)-(78) show regions where the spacetime is not dragged by the black hole rotation parameter. This means that the particles orbits would not be subject of the precession occurring due to the frame dragging effect body5 . Such a result is essentially due to the acoustic parameter that cancels the rotating effect.
@@ 7 Deflection angle
The metric (24) describes a slowly rotating acoustic black hole. We begin by rewriting such a metric in the equatorial plan in order to get the required the deflection angle of light. Indeed for $\theta=\frac{\pi}{2}$, we have
$ds^{2}=-\left[F(r)+2a\left(1-F(r)\right)\frac{d\phi}{dt}\right]dt^{2}+\frac{dr^{2}}{F(r)}+r^{2}d\phi^{2}.$ (79)
The equations given in (33) can be used to compute the factor $\frac{d\phi}{dt}$ in this equation. The expression for such a quantity as a function of the impact parameter $b=\frac{L}{E}$, can be given by
$\frac{d\phi}{dt}=\frac{F(r)+ba\left(1-F(r)\right)}{br^{2}-a\left(1-F(r)\right)}.$ (80)
@@
$dr_{*}$ $=\frac{dr}{\sqrt{F(r)\left(F(r)+2a\left(1-F(r)\right)\frac{d\phi}{dt}\right)}},$ (81)
$f(r_{*})$ $=\frac{r}{\sqrt{F(r)+2a\left(1-F(r)\right)\frac{d\phi}{dt}}},$ (82)
@@
$\iint_{D_{R}}K_{\mathcal{G}}dS+\oint_{\partial D_{R}}kdt+\sum n_{i}=2\pi\chi\left(D_{R}\right).$ (84)
@@
$k\left(\gamma_{R}\right)=\big|\nabla_{\gamma_{R}}\dot{\gamma_{R}}\big|.$ (85)
Given that $\gamma_{R}=R=cte$ is verified by the geodesic $\gamma_{R}$ def1 , where $R$ is the "geometrical size" of the region $D_{R}$, the radial component of $k\left(\gamma_{R}\right)$ becomes
$\left(\nabla_{\gamma_{R}}\dot{\gamma_{R}}\right)^{r}=\dot{\gamma_{R}}^{\phi}+\partial_{\phi}\dot{\gamma_{R}}^{r}+\Gamma^{r}_{\phi\phi}\left(\dot{\gamma_{R}}^{\phi}\right)^{2}.$ (86)
Following def1 , the second term provides
$\oint_{\partial D_{R}}kdt=\pi+\widehat{\alpha}$. (87)
@@
$\widehat{\alpha}=-\int_{0}^{\pi}\int_{\frac{b}{\sin\phi}}^{\infty}K_{\mathcal{G}}dS,$ (88)
@@
$K_{\mathcal{G}}=\frac{\mathbf{R}}{2},$ (89)
@@
$K_{\mathcal{G}}=\frac{18abM(\xi+1)}{r^{5}}-\frac{2M(\xi+1)}{r^{3}}.$ (90)
@@
$\widehat{\alpha}=\frac{4M(\xi+1)}{b}-\frac{4aM(\xi+1)}{b^{2}}+\mathcal{O}\left(M^{2},a^{2}\right),$ (91)
where $M$ higher orders are not included. At the limit $\xi\to 0$, the deflection angle (91) becomes
$\widehat{\alpha}\simeq\frac{4M}{b}\mp\frac{4aM}{b^{2}},$ (92)
where the $\mp$-branches would pertain to the pro-/retrograde motion. The obtained result matches precisely the LTBH black hole deflection of light C2 . To analyse the deflection of light by a slowly rotating acoustic black hole, we plot the associated behaviors as a function of the impact parameter $b$ in figure (7) for different values of the acoustic parameter $\xi$ and the spin $a$.
@@ Appendix A On the placement of critical radii
Further investigation could be carried regarding the LTABH who’s spacetime is described by the metric of equation (24). Indeed, we have seen that such a spacetime strongly depend on the value of the acoustic parameter $\xi$. While the horizon radius is always equal to $2M$, distinctions are also present. For convenience, we use the following notations
$r_{1}:=r_{ps},\,\,R_{s}\left(r_{1}\right)=R_{s}\left(r_{ps}\right):=R_{s},\quad r_{2}:=r_{as},\,\,R_{s}\left(r_{2}\right)=R_{s}\left(r_{as}\right):=R_{as}.$ (93)
@@ Appendix B The matrix $\mathcal{G}_{\mu\nu}$
$\mathcal{G}_{\mu\nu}=\begin{pmatrix}\mathcal{G}_{tt}&\mathcal{G}_{tr}&\mathcal{G}_{t\theta}&\mathcal{G}_{t\phi}\\ \mathcal{G}_{rt}&\mathcal{G}_{rr}&\mathcal{G}_{r\theta}&\mathcal{G}_{r\phi}\\ \mathcal{G}_{\theta t}&\mathcal{G}_{\theta r}&\mathcal{G}_{\theta\theta}&\mathcal{G}_{\theta\phi}\\ \mathcal{G}_{\phi t}&\mathcal{G}_{\phi r}&\mathcal{G}_{\phi\theta}&\mathcal{G}_{\phi\phi}\end{pmatrix}$, (94)
@@
$\mathcal{G}_{tt}$ $=\frac{c_{s}}{\sqrt{c_{s}^{2}-v_{\alpha}v^{\alpha}}}\,g_{tt}\left(c_{s}^{2}-v_{i}v^{i}\right),$ (95)
$\mathcal{G}_{tr}$ $=\frac{c_{s}}{\sqrt{c_{s}^{2}-v_{\alpha}v^{\alpha}}}\,\left(-v_{t}v_{r}+g_{tr}(c_{s}^{2}-v_{i}v^{i})\right),$ (96)
$\mathcal{G}_{t\theta}$ $=\frac{c_{s}}{\sqrt{c_{s}^{2}-v_{\alpha}v^{\alpha}}}\,\left(-v_{t}v_{\theta}+g_{t\theta}(c_{s}^{2}-v_{i}v^{i})\right),$ (97)
$\mathcal{G}_{t\phi}$ $=\frac{c_{s}}{\sqrt{c_{s}^{2}-v_{\alpha}v^{\alpha}}}\,\left(-v_{t}v_{\phi}+g_{t\phi}(c_{s}^{2}-v_{i}v^{i})\right),$ (98)
$\mathcal{G}_{rt}$ $=\frac{c_{s}}{\sqrt{c_{s}^{2}-v_{\alpha}v^{\alpha}}}\,\left(-v_{r}v_{t}+g_{rt}(c_{s}^{2}-v_{t}v^{t})\right),$ (99)
$\mathcal{G}_{rr}$ $=\frac{c_{s}}{\sqrt{c_{s}^{2}-v_{\alpha}v^{\alpha}}}\,\left(g_{rr}(c_{s}^{2}-v_{\alpha}v^{\alpha})+v_{r}v_{r}\right),$ (100)
$\mathcal{G}_{r\theta}$ $=\frac{c_{s}}{\sqrt{c_{s}^{2}-v_{\alpha}v^{\alpha}}}\,\left(g_{r\theta}(c_{s}^{2}-v_{\alpha}v^{\alpha})+v_{r}v_{\theta}\right),$ (101)
$\mathcal{G}_{r\phi}$ $=\frac{c_{s}}{\sqrt{c_{s}^{2}-v_{\alpha}v^{\alpha}}}\,\left(g_{r\phi}(c_{s}^{2}-v_{\alpha}v^{\alpha})+v_{r}v_{\phi}\right),$ (102)
$\mathcal{G}_{\theta t}$ $=\frac{c_{s}}{\sqrt{c_{s}^{2}-v_{\alpha}v^{\alpha}}}\,\left(-v_{\theta}v_{t}+g_{\theta t}(c_{s}^{2}-v_{t}v^{t})\right),$ (103)
$\mathcal{G}_{\theta r}$ $=\frac{c_{s}}{\sqrt{c_{s}^{2}-v_{\alpha}v^{\alpha}}}\,\left(g_{\theta r}(c_{s}^{2}-v_{\alpha}v^{\alpha})+v_{\theta}v_{r}\right),$ (104)
$\mathcal{G}_{\theta\theta}$ $=\frac{c_{s}}{\sqrt{c_{s}^{2}-v_{\alpha}v^{\alpha}}}\,\left(g_{\theta\theta}(c_{s}^{2}-v_{\alpha}v^{\alpha})+v_{\theta}v_{\theta}\right),$ (105)
$\mathcal{G}_{\theta\phi}$ $=\frac{c_{s}}{\sqrt{c_{s}^{2}-v_{\alpha}v^{\alpha}}}\,\left(g_{\theta\phi}(c_{s}^{2}-v_{\alpha}v^{\alpha})+v_{\theta}v_{\phi}\right),$ (106)
$\mathcal{G}_{\phi t}$ $=\frac{c_{s}}{\sqrt{c_{s}^{2}-v_{\alpha}v^{\alpha}}}\,\left(-v_{\phi}v_{t}+g_{\phi t}(c_{s}^{2}-v_{t}v^{t})\right),$ (107)
$\mathcal{G}_{\phi r}$ $=\frac{c_{s}}{\sqrt{c_{s}^{2}-v_{\alpha}v^{\alpha}}}\,\left(g_{\phi r}(c_{s}^{2}-v_{\alpha}v^{\alpha})+v_{\phi}v_{r}\right),$ (108)
$\mathcal{G}_{\phi\theta}$ $=\frac{c_{s}}{\sqrt{c_{s}^{2}-v_{\alpha}v^{\alpha}}}\,\left(g_{\phi\theta}(c_{s}^{2}-v_{\alpha}v^{\alpha})+v_{\phi}v_{\theta}\right),$ (109)
$\mathcal{G}_{\phi\phi}$ $=\frac{c_{s}}{\sqrt{c_{s}^{2}-v_{\alpha}v^{\alpha}}}\,\left(g_{\phi\phi}(c_{s}^{2}-v_{\alpha}v^{\alpha})+v_{\phi}v_{\phi}\right).$ (110)